Given two independent random samples with the following results:

n1=128  n2= 253 x1=58  x2=73

Can it be concluded that there is a difference between the two population proportions? Use a significance level of α=0.01 for the test.

Step 1 of 6: State the null and alternative hypotheses for the test.

Step 2 of 6: Find the values of the two sample proportions, pˆ1 and pˆ2. Round your answers to three decimal places.

Step 3 of 6: Compute the weighted estimate of p, p‾‾p‾. Round your answer to three decimal places.

Step 4 of 6: Compute the value of the test statistic. Round your answer to two decimal places.

Step 5 of 6: Determine the decision rule for rejecting the null hypothesis H0H0. Round the numerical portion of your answer to three decimal places.

Step 6 of 6: Make the decision for the hypothesis test.
Answer #1

Below are the null and alternative Hypothesis,
Null Hypothesis, H0: p1 = p2
Alternate Hypothesis, Ha: p1 ≠ p2

p1cap = X1/N1 = 58/128 = 0.453
p1cap = X2/N2 = 73/253 = 0.289

pcap = (X1 + X2)/(N1 + N2) = (58+73)/(128+253) = 0.344


Test statistic
z = (p1cap - p2cap)/sqrt(pcap * (1-pcap) * (1/N1 + 1/N2))
z = (0.453-0.289)/sqrt(0.344*(1-0.344)*(1/128 + 1/253))
z = 3.18


Rejection Region
This is two tailed test, for α = 0.01
Critical value of z are -2.58 and 2.58.
Hence reject H0 if z < -2.58 or z > 2.58

Reject the null hypothesis